# Global identification of dynamic panel models with interactive effects

### v1

@@ 6 Concluding remarks
In this paper, we demonstrate that the parameters of a dynamic panel data model with interactive effects can be globally identified. We extend the seminal result of Anderson and Rubin (1956) to dynamic panels with interactive effects, a problem that was previously thought to be non-globally identifiable.
@@
Our proof shows that global identification is achievable when $T\geq 2(\bar{r}+1)$. An interesting open question is whether global identification can be attained under weaker conditions, such as when $T$ satisfies the Lebdermann bound plus one. The requirement for an additional period, beyond the Lebdermann bound, arises from the dynamic nature of the model, which includes an extra parameter compared to the static case.
Appendix A Explicit expression for $\mathbf{O}_{i,j}$ in Section 4 and $\bar{r}=1$
As $\mathbf{O}$ is symmetric, then $\mathbf{O}_{i,j}=\mathbf{O}_{j,i}$.The positions of $\mathbf{O}$ used in section 4.3 are:
$\mathbf{O}_{1,2}$ $=\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\Psi$
$\mathbf{O}_{1,3}$ $=f_{3}\,\Psi-\left(d_{1}+\Psi\right)\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
$\mathbf{O}_{1,4}$ $=f_{4}\,\Psi-\left(\alpha^{2}\,\widetilde{\alpha}-\alpha^{3}\right)\,\left(d_{1}+\Psi\right)+f_{3}\,\Psi_{1}\,\left(\alpha-\widetilde{\alpha}\right)$
@@ Appendix A Explicit expression for $\mathbf{O}_{i,j}$ in Section 4 and $\bar{r}=1$
$\mathbf{O}_{2,3}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(\Psi\,{f_{2}}^{2}+\Psi\,\left(\alpha-\widetilde{\alpha}\right)\,f_{2}+d_{2}\right)$
@@
$\quad+f_{2}\,f_{3}\,\Psi+f_{3}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
$\mathbf{O}_{2,4}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(f_{2}\,f_{3}\,\Psi+f_{3}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)\right)$
@@
$\quad-\left(\alpha^{2}\,\widetilde{\alpha}-\alpha^{3}\right)\,\left(f_{2}\,\Psi+\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)\right)$
$\quad+f_{2}\,f_{4}\,\Psi+f_{4}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
$\mathbf{O}_{3,4}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(d_{3}+{f_{3}}^{2}\,\Psi-f_{3}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,f_{3}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)\right)$
@@
$\quad-\left(\alpha^{2}\,\widetilde{\alpha}-\alpha^{3}\right)\,\left(f_{3}\,\Psi-\left(d_{1}+\Psi\right)\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)\right)$
$\quad+f_{3}\,f_{4}\,\Psi-f_{4}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,f_{4}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$

### v2

@@ 6 Concluding remarks
In this paper, we demonstrate that the parameters of a dynamic panel data model with interactive effects can be globally identified. We extend the seminal result of Anderson and Rubin, (1956) to dynamic panels with interactive effects, a problem that was previously thought to be non-globally identifiable.
@@
Our proof shows that global identification is achievable when $T\geq 2(\bar{r}+1)$. An interesting open question is whether global identification can be attained under weaker conditions, such as when $T$ satisfies the Ledermann bound plus one. The requirement for an additional period, beyond the Ledermann bound, arises from the dynamic nature of the model, which includes an extra parameter compared to the static case.
Appendix A Explicit expression for $O_{i,j}$ in Section 4 and $\bar{r}=1$
As $\mathbf{O}$ is symmetric, then $O_{i,j}=O_{j,i}$.The positions of $\mathbf{O}$ used in section 4.3 are:
$O_{1,2}$ $=\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\Psi$
[…]
@@ Appendix A Explicit expression for $O_{i,j}$ in Section 4 and $\bar{r}=1$
$O_{2,3}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(\Psi\,{f_{2}}^{2}+\Psi\,\left(\alpha-\widetilde{\alpha}\right)\,f_{2}+d_{2}\right)$
@@
[…]
@@
[…]
@@
[…]
@@ Appendix B Rotational indeterminacy in the presence of individual fixed-effects and $\bar{r}=2$
Appendix C Number of unique tridiagonal exclusion minors
Proposition 3.
If $T=2(\bar{r}+1)+1$, then there exists exactly one unique tridiagonal exclusion minor of dimension $\bar{r}+1$ of $\mathbf{\dot{O}}$.
Proof.
A $(\bar{r}+1)\times(\bar{r}+1)$ tridiagonal exclusion minor is defined by two index sets
$R=\{r_{1},\dots,r_{\bar{r}+1}\},\qquad C=\{c_{1},\dots,c_{\bar{r}+1}\},$
with $|R|=|C|=\bar{r}+1$, such that every selected entry lies outside the tridiagonal band, i.e.
$|r-c|\geq 2\quad\text{for all }r\in R,\ c\in C.$
For a fixed $R$, define the tridiagonal neighborhood
$N(R):=\{\,r-1,\,r,\,r+1:r\in R\,\}\cap[T].$
Then, any valid column set must be contained in the complement $T]\setminus N(R)$, so the number of admissible columns equals
$|[T]\setminus N(R)|=T-|N(R)|.$
In order to choose a $(\bar{r}+1)\times(\bar{r}+1)$ tridiagonal exclusion minor we require
$T-|N(R)|\;\geq\;\bar{r}+1.$
Let $|R|=\bar{r}+1$. Each $r\in R$ contributes up to three elements $\{r-1,r,r+1\}$ to $N(R)$. Overlaps may occur. We can check that
$|N(R)|\;\geq\;\bar{r}+2,$
with equality if and only if $R$ is a block of $\bar{r}+1$ consecutive integers that touches a boundary (i.e. $R=\{1,\dots,\bar{r}+1\}$ or $R=\{T-\bar{r},\dots,T\}$). Any other placement yields $|N(R)|>\bar{r}+2$.
Now, let $T=2(\bar{r}+1)+1$. In this case
$T-|N(R)|\;\leq\;2(\bar{r}+1)+1-(\bar{r}+2)=\bar{r}+1.$
Hence the inequality $T-|N(R)|\geq\bar{r}+1$ forces equality, so we must have $|N(R)|=\bar{r}+2$. By Step 1 this occurs only when
$R=\{1,\dots,\bar{r}+1\}\quad\text{or}\quad R=\{T-\bar{r},\dots,T\}.$
We can use this result to pin down the possible composition of $C$. If $R=\{1,\dots,\bar{r}+1\}$ then
$N(R)=\{1,2,\dots,\bar{r}+2\},\qquad[T]\setminus N(R)=\{\bar{r}+3,\dots,2(\bar{r}+1)+1\}.$
Thus $C$ is forced to be $\{\bar{r}+3,\dots,2(\bar{r}+1)+1\}$. If instead $R=\{T-\bar{r},\dots,T\}$ then symmetrically $C=\{1,\dots,\bar{r}+1\}$.
The two ordered pairs $(R,C)$ obtained above are transposes of each other. In a symmetric matrix they yield the same minor value. Therefore, up to equality, there is exactly one tridiagonal exclusion minor of dimension $\bar{r}+1$. $\square$